HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025, Page 6 of 7  Not logged in ELOG logo
ID Date Author Category Subject Text Attachments
  35   Tue Apr 1 11:47:14 2025 Jelena B, SiddharthGeneralLassie 2025-04-01_11-46-35-517_tcl1064.png    2025-04-01_11-46-35-517_tcl1064.png 
  34   Tue Apr 1 10:34:20 2025 Jelena B, SiddharthGeneralResolved issue with SIS18Beam is back with some fluctuations.
We did not have beam from 9:34 until 10:08,
around 35 minutes. 
 2025-04-01_10-33-38-948_tcl1064.png 
  33   Tue Apr 1 10:03:56 2025 Jelena B, SiddharthGeneralTechnical issue with SIS18 From HKR info: there is a technical issue
with SIS18 and we will have no beam for next
25 minutes. 
 2025-04-01_10-02-16-454_tcl1064.png 
  32   Tue Apr 1 08:36:25 2025 Jelena B, SiddharthGeneralLassie 2025-04-01_08-37-32-155_tcl1064.png   2025-04-01_08-37-32-155_tcl1064.png 
  31   Tue Apr 1 06:53:51 2025 Jelena B,SiddharthGeneralLassie Monitor   2025-04-01_06-49-59-989_tcl1064.png 
  30   Tue Apr 1 06:31:46 2025 MichaelGeneralLassie Monitor 2025-04-01_06-31-10-211_tcl1064.png   2025-04-01_06-31-10-211_tcl1064.png 
  29   Tue Apr 1 05:58:51 2025 MichaelGeneralLassie Monitor 2025-04-01_05-58-09-122_tcl1064.png   2025-04-01_05-58-09-122_tcl1064.png 
  28   Tue Apr 1 05:29:45 2025 MichaelGeneralLassie Monitor 2025-04-01_05-29-01-845_tcl1064.png   2025-04-01_05-29-01-845_tcl1064.png 
  27   Tue Apr 1 04:59:14 2025 MichaelGeneralLassie Monitor 2025-04-01_04-58-31-351_tcl1064.png   2025-04-01_04-58-31-351_tcl1064.png 
  26   Tue Apr 1 04:27:10 2025 MichaelGeneralLassie Monitor 2025-04-01_04-25-53-171_tcl1064.png   2025-04-01_04-25-53-171_tcl1064.png 
  25   Tue Apr 1 03:55:59 2025 MichaelGeneralLassie Monitor 2025-04-01_03-55-21-183_tcl1064.png   2025-04-01_03-55-21-183_tcl1064.png 
  24   Tue Apr 1 03:52:56 2025 Shahab, Michael, Yue, Zac, CarloDAQgood filesall settings are done.
any files coming after this one is good:
  
  23   Tue Apr 1 03:50:12 2025 Shahab, Yue, Carlo, Zac, Esther, MichaelDAQTuners of the both cavities set410 MHz cavity resonance frequency tuned.
motors both on 50 mm
245 MHz cavity resonance frequency tuned.
  
  22   Tue Apr 1 03:27:44 2025 MichaelGeneralLassie Monitor 2025-04-01_03-22-05-395_tcl1064.png    2025-04-01_03-22-05-395_tcl1064.png 
  21   Tue Apr 1 02:55:06 2025 Shahab, Michael, Yue, Zac, CarloDAQnew RSA recordinganything from RSA01-2025.04.01.02.54.25.976.tiq
onwards is good data
 IMG_4955.JPGIMG_4956.JPGIMG_4954.JPG 
  20   Tue Apr 1 02:37:28 2025 H. Weick, S. LitvinovAcceleratorIsoc-setting +scrapersFirst a setting with orbit bump for injection
on central orbit was tried.
- quadrupoles tuned and scrapers optimized
 IMG_20250401_013626.jpgIMG_20250401_013636.jpgIMG_20250401_013646.jpgIMG_20250401_023358.jpg 
  19   Tue Apr 1 02:20:36 2025 ruijiuAnalysischange back the frequency of 194Pb82frequency of 194Pb82+ at 211 408606000
frequency of 194Pb82+ at 126 244001000
 Screenshot_2025-04-01_02-20-09.png 
  18   Tue Apr 1 02:03:33 2025 ruijiuAnalysisThe frequency of 194Pb82+ has been changed from 408.522 MHz to 408.637 MHzthis is the new setting after we lost the
old one.
 Screenshot_2025-04-01_02-03-14.png 
  17   Tue Apr 1 02:01:06 2025 Zac, Shahab, Ruijiu, Yuri, Sergey, Wolfram, Carlo, MichaelAcceleratorSettings lostSettings lost after the beam was lost. Trying
to recover particle identification now that
beam has returned.
  
  16   Mon Mar 31 23:34:58 2025 ruijiu, shahab, yuri, wolfram, sergeyAnalysisidentificaton of 194Pb82+datafile = "/data.local/G22-00018_00203/results/analyzers/RSA01/RSA01-2025.03.31.22.18.16.965.tiq_spectrum.npz"
filep = "/data.local/G22-00018_00203/scripts/OnlineDataAnalysisSystem/rionid_exercises/exercise_194Pb_Exp_2025/New_TEline2022-ESR-194Pb-208Pb-nostripper_524MeV.lpp"
alphap = "1.4"
 8x 
ELOG V3.1.5-fc6679b